# Magnetic Penrose process in the magnetized Kerr spacetime

### 24 Jan 2024 v1

@@ Abstract
It is well-established that the magnetic Penrose process (MPP) could be highly efficient (efficiency can even exceed $100\%$) for extracting the energy from a Kerr black hole, if it is immersed in a mG order magnetic field. Considering the exact solution of the magnetized Kerr spacetime, here we derive the exact expression of efficiency ($\eta_{\rm MPP}$) for MPP, which is valid for both the Kerr black hole (BH) as well as Kerr superspinar (SS), and also from the weak magnetic field to an ultra-strong magnetic field $(B)$ which can even distort the original Kerr geometry. We show that although the value of $\eta_{\rm MPP}$ increases upto a certain value of ultra-strong magnetic field ($B_{p}$), it decreases to zero for $B>B_{p}$, in case of the Kerr BHs. On the other hand, $\eta_{\rm MPP}$ shows the opposite behavior in case of the Kerr SSs. One intriguing feature that emerges is, $\eta_{\rm MPP}$ acquires the maximum value for the Kerr parameter $a_{*}\approx 0.786$ (unlike $a_{*}=1$ for the ordinary PP), decreases for the range $0.786<a_{*}\leq 1$, and reaches to $20.7\%$ for $a_{*}=1$ with a few limitations. This indicates that the BH starts to expel the effect of magnetic field for $a_{*}>0.786$, and is fully expelled from the extremal Kerr BH due to the gravitational Meissner effect. As a special case of MPP, we also study the ordinary Penrose process (PP) for the magnetized Kerr spacetime. We show that the efficiency of PP decreases with increasing the magnetic field for the Kerr BH. In case of the Kerr SS, the efficiency of PP decreases from $10^{3}\%$ to 00 for increasing the value of magnetic field from 00 to a specific value of ultra-strong magnetic field. Thus, the MPP for Kerr BHs, Kerr SSs and the ordinary PP for Kerr SSs can be superefficient for the astrophysical applications to powering engines in the high-energy sources like active galactic nuclei and quasars, in the weak magnetic fields. Secondly, the origin of ultrahigh-energy ($\gtrsim 10^{20}$ eV) cosmic rays received on Earth could also be explained by the MPP mechanism, as $\eta_{\rm MPP}$ could reach maximum $\sim 10^{22}\%$ for a magnetized Kerr BH. It is almost impossible to extract the energy from a BH (SS) through MPP (PP) in the ultra-strong magnetic fields.
@@ I Introduction
In 1985, the magnetic version of the PP was proposed by Wagh et al. Wagh85 , which is known as the magnetic Penrose process (MPP). This is almost similar to the PP with two major modifications. First one is, the Kerr BH should be surrounded by the magnetic fields. Second one is, the neutral particle should be split into two charged fragments. One charged particle attains the negative energy, and falls into the BH. Another charged particle escapes to infinity with energy exceeding the energy of the incident particle to conserve the total energy Wagh85 . Later, it was shown bhat ; ps86 that the efficiency of MPP could even exceed $100\%$ if a stellar mass Kerr BH is surrounded by a mG order magnetic fields dh . Note that the efficiency of MPP does not only depend on the geometry of the spacetime, but it also depends on the non-vanishing components of the four-potential $A_{\mu}$ of the electromagnetic fields and some other parameters. Remarkably, the condition of the relative velocity between the two fragments $c/2$ (where $c$ is the speed of light in vacuum) for PP bpt ; Wald74a ; Wald74b , is nicely circumvented in MPP. This is because the required energy could now come from the electromagnetic fields leaving the relative velocity between the two fragments to be free dh . However, the formalism of MPP was developed Wagh85 ; ps86 ; bhat based on the ordinary Kerr BH assuming the weak magnetic field, and the components of $A_{\mu}$ was presumed to be the linear order in $B$ dh .
@@ II Energy extraction through the magnetic Penrose Process from a stationary and axisymmetric spacetime
A stationary and axisymmetric spacetime is generally expressed as
@@ IV.1 Efficiency of magnetic Penrose process for the magnetized Kerr black hole
Figure. 1: $\eta_{\rm MPP}$ versus $B$ for the Kerr BHs surrounded by an ultra-strong magnetic field. It shows that the peak ($\eta_{\rm max}$) of $\eta_{\rm MPP}$ starts to decrease for $a_{*}>0.79$, and $\eta_{\rm MPP}\leq 20.7\%$ for $a_{*}=1$, as seen from FIG. 4. See Sec. IV.1 for details.
Figure. 1: $\eta_{\rm MPP}$ versus $B$ for the Kerr BHs surrounded by an ultra-strong magnetic field. It shows that the peak ($\eta_{\rm max}$) of $\eta_{\rm MPP}$ starts to decrease for $a_{*}>0.79$, and $\eta_{\rm MPP}\leq 20.7\%$ for $a_{*}=1$, as seen from FIG. 4. See Sec. IV.1 for details.
@@
Figure. 3: $\eta_{\rm MPP}^{\rm exBH}$ versus $B$ for an extremal Kerr BH surrounded by a magnetic field of $B\sim 10^{-7}M^{-1}$. It shows that the efficiency for an extremal Kerr BH remains constant (as seen from the solid red line of FIG. 2), then decreases and vanishes in the comparatively weak magnetic field shown in FIG. 2. See Sec. IV.1 for details.
Figure. 3: $\eta_{\rm MPP}^{\rm exBH}$ versus $B$ for an extremal Kerr BH surrounded by a magnetic field of $B\sim 10^{-7}M^{-1}$. It shows that the efficiency for an extremal Kerr BH remains constant (as seen from the solid red line of FIG. 2), then decreases and vanishes in the comparatively weak magnetic field shown in FIG. 2. See Sec. IV.1 for details.
@@ V Summary and discussion
Although the upper limit of the magnetic fields in the X-ray corona of the BH Cygnus X-1 rises to $10^{7}$ G santo , emission models of Sagittarius A* and M87* provide $B\sim 30-100$ G eatough and $B\sim 1-30$ G eht7 respectively, the ultra-strong magnetic fields ($\sim$ Eq. 1) around a BH has not been detected yet. It is still unclear whether such a ultra-strong magnetic fields ($B>0.36M^{-1}$) exist in the Universe. If the ultra-strong magnetic field even exists around a collapsed object, it may not be detectable due to the magnetic shielding shld . However, our result could be verified if such an ultra-strong magnetic field is discovered around a collapsed object in future.
@@
The production mechanisms of ultrahigh-energy (e.g. EeV order) cosmic rays which we receive on Earth from the outer space, remains unclear. Almost all the currently available mechanisms are based on the electromagnetic interaction between the accelerated charged particles. Intriguingly, the highest energy of a detected cosmic ray is $3\times 10^{20}$ eV nasa . Our result shows that $\eta_{\rm MPP}$ could reach maximum $\sim 3\times 10^{22}\%$ (see FIG. 2) in case of a magnetized Kerr BH and much more for a magnetized Kerr SS. In other words, FIG. 2 shows that a charged particle of $1$ eV could escape to infinity with the energy $10^{20}$ eV from a magnetized Kerr BH of $a_{*}\sim 0.79$, which is surrounded by $B\sim 0.2M^{-1}$. Therefore, it could easily explain the production mechanism of $100$ EeV ($\sim 10^{20}$ eV) cosmic rays. As the number $3\times 10^{22}\%$ is independent of the mass of the BH, this huge energy could be originated from a supermassive or a stellar mass Kerr BH if it is immersed in the (ultra-)strong magnetic fields (see FIG. 2) that could distort the original Kerr spacetime ag2 . Therefore, this specific observational implication is unique in the sense that it arises due to the presence of ultra-strong magnetic fields (e.g., $B\sim 0.2M^{-1}$) around a Kerr BH, for which the effect of $B$ needs to be introduced to the original Kerr metric. This ultrahigh-energy production, in fact, cannot be explained by the ordinary Penrose process originated from a (magnetized) Kerr BH or by the BZ mechanism, as the maximum efficiency of Penrose process for a Kerr BH reaches only upto $20.7\%$.
@@
Acknowledgements : One of us (CC) thanks M. Astorino for useful discussions regarding the magnetized Kerr spacetime, and S. V. Dhurandhar for an insightful discussion on the magnetic Penrose process. CC also thanks the Inter-University Centre for Astronomy and Astrophysics (IUCAA), Pune, India, as the part of this work was done during the academic visit to IUCAA under the visiting associateship programme.

### 04 Mar 2024 v2

@@ Abstract
It is well-established that the magnetic Penrose process (MPP) could be highly efficient (efficiency can even exceed $100\%$) for extracting the energy from a Kerr black hole, if it is immersed in a mG order magnetic field. Considering the exact solution of the magnetized Kerr spacetime, here we derive the exact expression of efficiency ($\eta_{\rm MPP}$) for MPP, which is valid for both the Kerr black hole (BH) as well as Kerr superspinar (SS), and also from the weak magnetic field to an ultra-strong magnetic field $(B)$ which can even distort the original Kerr geometry. We show that although the value of $\eta_{\rm MPP}$ increases upto a certain value of ultra-strong magnetic field ($B_{p}$), it decreases to zero for $B>B_{p}$, in case of the Kerr BHs. On the other hand, $\eta_{\rm MPP}$ shows the opposite behavior in case of the Kerr SSs. One intriguing feature that emerges is, $\eta_{\rm MPP}$ acquires the maximum value for the Kerr parameter $a_{*}\approx 0.786$ (unlike $a_{*}=1$ for the ordinary PP), decreases for the range $0.786<a_{*}\leq 1$, and reaches to $20.7\%$ for $a_{*}=1$ with a few limitations. This indicates that the BH starts to expel the effect of magnetic field for $a_{*}>0.786$, and is fully expelled from the extremal Kerr BH due to the gravitational Meissner effect. As a special case of MPP, we also study the ordinary Penrose process (PP) for the magnetized Kerr spacetime. We show that the efficiency of PP decreases with increasing the magnetic field for the Kerr BH. In case of the Kerr SS, the efficiency of PP decreases from $10^{3}\%$ to 00 for increasing the value of magnetic field from 00 to a specific value of ultra-strong magnetic field. Thus, the MPP for Kerr BHs, Kerr SSs and the ordinary PP for Kerr SSs can be superefficient for the astrophysical applications to powering engines in the high-energy sources like active galactic nuclei and quasars, in the weak magnetic fields. Our strong magnetic field result of MPP could be important to the primordial BHs in the early Universe immersed in the primordial magnetic fields, and to the transmuted BHs which are formed by collapsing and/or by merging of the magnetized progenitors. It is almost impossible to extract the energy from a BH (SS) through MPP (PP) in the ultra-strong magnetic fields.
@@ I Introduction
In 1985, the magnetic version of the PP was proposed by Wagh et al. Wagh85 , which is known as the magnetic Penrose process (MPP). This is almost similar to the PP with two major modifications. First one is, the Kerr BH should be surrounded by the magnetic fields. Second one is, the neutral particle should be split into two charged fragments. One charged particle attains the negative energy, and falls into the BH. Another charged particle escapes to infinity with energy exceeding the energy of the incident particle to conserve the total energy Wagh85 . Later, it was shown bhat ; ps86 that the efficiency of MPP could even exceed $100\%$ if a stellar mass Kerr BH is surrounded by a mG order magnetic fields dh . Note that the efficiency of MPP does not only depend on the geometry of the spacetime, but it also depends on the non-vanishing components of the four-potential $A_{\mu}$ of the electromagnetic fields and some other parameters. Remarkably, the condition of the relative velocity between the two fragments $c/2$ (where $c$ is the speed of light in vacuum) for PP bpt ; Wald74a ; Wald74b , is nicely circumvented in MPP. This is because the required energy could now come from the electromagnetic fields leaving the relative velocity between the two fragments to be free dh . However, the formalism of MPP was developed Wagh85 ; ps86 ; bhat based on the ordinary Kerr BH assuming the weak magnetic field, and the components of $A_{\mu}$ was presumed to be the linear order in $B$ dh . In another work, it is recently proposed dyson that the accretion can happen in a superradiant manner, and the angular momentum can be extracted from the dyonic Kerr-Newman BH via Penrose process in an enhanced rate due to the dipolar electric field created by the rotating magnetic charge of the BH.
@@ II Energy extraction through the magnetic Penrose Process from a stationary and axisymmetric spacetime
We consider a stationary and axisymmetric spacetime with the line element
@@ IV.1 Efficiency of magnetic Penrose process for the magnetized Kerr black hole
Figure. 1: $\eta_{\rm MPP}$ versus $B$ for the Kerr BHs surrounded by an ultra-strong magnetic field. It shows that the peak ($\eta_{\rm max}$) of $\eta_{\rm MPP}$ starts to decrease for $a_{*}>0.79$, and $\eta_{\rm MPP}\leq 20.7\%$ for $a_{*}=1$, as seen from FIG. 4. As the value of $\eta_{\rm MPP}$ is very low for $a_{*}=1$ compared to the values shown along the Y-axis in the present plot, we draw FIG. 4 separately. See Sec. IV.1 for details.
Figure. 1: $\eta_{\rm MPP}$ versus $B$ for the Kerr BHs surrounded by an ultra-strong magnetic field. It shows that the peak ($\eta_{\rm max}$) of $\eta_{\rm MPP}$ starts to decrease for $a_{*}>0.79$, and $\eta_{\rm MPP}\leq 20.7\%$ for $a_{*}=1$, as seen from FIG. 4. As the value of $\eta_{\rm MPP}$ is very low for $a_{*}=1$ compared to the values shown along the Y-axis in the present plot, we draw FIG. 4 separately. See Sec. IV.1 for details.
@@
Figure. 3: $\eta_{\rm MPP}^{\rm exBH}$ versus $B$ for an extremal Kerr BH ($a_{*}=1$) surrounded by a magnetic field of $B\sim 10^{-7}M^{-1}$. It shows that the efficiency for an extremal Kerr BH remains constant (as seen from the solid red line of FIG. 2), then decreases and vanishes in the comparatively stronger magnetic field shown in FIG. 2. See Sec. IV.1 for details.
Figure. 3: $\eta_{\rm MPP}^{\rm exBH}$ versus $B$ for an extremal Kerr BH ($a_{*}=1$) surrounded by a magnetic field of $B\sim 10^{-7}M^{-1}$. It shows that the efficiency for an extremal Kerr BH remains constant (as seen from the solid red line of FIG. 2), then decreases and vanishes in the comparatively stronger magnetic field shown in FIG. 2. See Sec. IV.1 for details.
@@ V Summary and discussion
Although the upper limit of the magnetic fields in the X-ray corona of the BH Cygnus X-1 rises to $10^{7}$ G santo , emission models of Sagittarius A* and M87* provide $B\sim 30-100$ G eatough and $B\sim 1-30$ G eht7 respectively, the ultra-strong magnetic fields ($\sim$ Eq. 1) around a BH has not been detected yet. It is still unclear whether such a ultra-strong magnetic fields ($B>0.36M^{-1}$) exist in the Universe. If the ultra-strong magnetic field even exists around a collapsed object, it may not be detectable due to the magnetic shielding shld . However, our result could be verified if such a ultra-strong magnetic field is discovered around a collapsed object in future.
@@
The production mechanisms of ultrahigh-energy (e.g. EeV order) cosmic rays which we receive on Earth from the outer space, remains unclear. Almost all the currently available mechanisms are based on the electromagnetic interaction between the accelerated charged particles. Intriguingly, the highest energy of a detected cosmic ray is $3\times 10^{20}$ eV nasa . Our result shows that $\eta_{\rm MPP}$ could reach maximum upto $\sim 3\times 10^{22}\%$ (note that the number is independent of the mass of the BH; see FIG. 2) in case of a magnetized Kerr BH and much more for a magnetized Kerr SS. For example, FIG. 2 shows that a charged particle of $1$ eV could escape to infinity with the energy $10^{20}$ eV from a magnetized Kerr BH of $a_{*}\sim 0.79$, which is surrounded by $B\sim 0.2M^{-1}$. Thus, the existence of any ultra-strong magnetized BHs in reality  [^4] It is not realistic to expect such a high magnetic field ($B\sim 0.2M^{-1}$) around a SMBH as of now, because $B\sim 0.2M^{-1}$ is equivalent to $10^{10}$ G for a SMBH of $M\sim 10^{9}M_{\odot}$, and such a strong magnetic field around a SMBH has not been detected yet. could have explained the production mechanism of EeV range energy cosmic rays.
A ultra-strong magnetic field in the order of $10^{20}$ G may be originated (see gru ; grasso and references therein) in the early Universe, and the primordial black holes (PBHs) could be immersed in that magnetic fields. Our exact result of MPP in the presence of ultra-strong magnetic fields could be applicable to those strongly magnetized PBHs. Our result could also be important to those BHs which are formed by collapse induced  [^5] These BHs are generally known as the transmuted BHs dasg . If a BH is formed by collapse induced by an endoparasitic BH through capturing of dark matter in a magnetar of $M\sim 2M_{\odot}$, one may expect a comparable strong magnetic field ($B\sim 0.2M^{-1}$ equivalent to $10^{18}$ G in this case) around that BH, or the BH formed by merging of several such strongly magnetized progenitors. Note that such a strongly-magnetized BH and/or a near-solar mass BH has not been detected yet (but, see dasg ; gw1 ; gw2 ). by an endoparasitic BH through capturing of dark matter into a neutron star gold ; dasg (or, white dwarf cbns , magnetar etc.) with magnetic field, and by merging of a BH with one (or multiple) magnetized neutron star(s) or magnetar(s) lyu83 ; east ; shld . In such cases, the magnetic field could be prevented from sliding off the newly-formed BH as shown in lyu ; lyu83 (but see brp ), and one can calculate $\eta_{\rm MPP}$ of such magnetized BHs by using our exact expression of MPP.
@@
Acknowledgements : We thank the referees for constructive comments that helped to improve the manuscript from the observational point of view. One of us (CC) thanks M. Astorino for useful discussions regarding the magnetized Kerr spacetime. CC also thanks S. V. Dhurandhar for an useful discussion on the magnetic Penrose process. CC also thanks the Inter-University Centre for Astronomy and Astrophysics (IUCAA), Pune, India, as the part of this work was done during the academic visit to IUCAA under the visiting associateship programme. CC dedicates this paper to his mother, Shree Snigdha Chakraborty who was his constant inspiration in all fields of life, and unfortunately passed away on September 11, 2023.